OpenStudy (anonymous):

Find the General Solution of \[\sin^2 x + \sin^2 2x = \sin^2 3x\]?

OpenStudy (anonymous):

@shubhamsrg @satellite73 @RadEn

OpenStudy (shubhamsrg):

apply there identities, you'll get sin^2 x + 4 sin^2 x cos^2 x = (-4sin^3 x + 3sinx)^2 = sin^2 x( 3 - 4sin^2 x)^2 one clear solution will be sin^2 x=0 or sin x =0 ..generalize this we are left with 1 + 4 cos^2 x = 9 + 16 sin^4 x - 24 sin^2 x =>1 + 4 - 4sin^2 x = 9 + 16 sin^4 x - 24 sin^2x => 16sin^4 x - 20 sin^2 x + 4 =0 => 4 sin^4 x - 5 sin^2 x + 4 =0 its a quadratic eqn in sin^2 x which can be easily solved (5= 4+1) rest should be easier..

OpenStudy (raden):

sin^2 x = sin^2 x :) sin^ 2x = (2sinxcosx)^2 = 4sin^2 x cos^2 x = 4sin^2 x - 4sin^4 x (sin3x)^2 = (-4sin^3 x + 3sinx)^2 = 16sin^6 x - 24sin^4 x + 9sin^2 x now, u get the equation in sinx form... rearreange giving us : 16sin^6 x -20sin^4 x + 4sin^2 x = 0 or 4sin^6 x -5sin^4 x + sin^2 x = 0 sin^2 x (4sin^4 x - 5sin^2 x + 1) = 0 sin^2 x (4sin^2 x - 1)(sin^2 x - 1) = 0 sin^2 x (2sinx+1)(2sinx-1)(sinx+1)(sinx-1) = 0 now, it looks be beatifull form right ?